# Transforming Image Super-Resolution: A ConvFormer-based Efficient Approach

### v1

@@ V Conclusion
The field of Single-Image Super-Resolution (SISR) has undergone significant advancements in recent years, largely attributed to the deep learning methods. In this paper we propose a Transformer-like, convolutional network for lightweight SR tasks, named CFSR, which achieves an advanced performance. CFSR employs a large kernel feature mixer (LK Mixer) as an efficient substitute for the complicate self-attention module, enabling the modeling of extensive receptive fields while significantly reducing computational overhead. Moreover, we propose an edge-preserving feed-forward network (EFN) to extract local features while preserving high-frequency information. To enhance the edge-preserving capabilities of our network, we propose the edge-preserving depth-wise convolution (EDC), which provides richer high-frequency information without introducing extra complexity during inference by a re-parameterizing strategy. To understand the influence of these components, detailed ablation studies are provided. Comprehensive experiments highlight the superior performance of CFSR, showcasing its ability to outclass existing state-of-the-art methods while maintaining a lean profile in terms of parameter count and computational complexity. Overall, our proposed method holds great potential for advancing the field of SISR and enabling the deployment of super-resolution algorithms on resource-constrained devices.

### v2

@@ V Conclusion
In recent years, the field of Single-Image Super-Resolution (SISR) has witnessed significant advancements, largely due to the adoption of deep learning techniques. In this paper, we have proposed a transformer-like convolutional network for lightweight super-resolution tasks, named CFSR, which achieves state-of-the-art performance. CFSR utilizes a large kernel feature mixer (LK Mixer) as an efficient alternative to the computationally intensive self-attention module, effectively modeling extensive receptive fields while substantially reducing computational overhead. Additionally, we have introduced an edge-preserving feed-forward network (EFN) to extract local features while preserving high-frequency information. To further enhance the edge-preserving capabilities of our network, we proposed the edge-preserving depth-wise convolution (EDC), which enriches high-frequency information without adding extra computational complexity during inference through a re-parameterization strategy. We conducted detailed ablation studies to understand the influence of these components. Comprehensive experiments demonstrate the superior performance of CFSR, showcasing its ability to surpass existing state-of-the-art methods while maintaining a lean profile in terms of parameter count and computational complexity. Overall, our proposed method holds significant potential for advancing the field of SISR and facilitating the deployment of super-resolution algorithms on resource-constrained devices.